OpenStudy (anonymous):

find the value of \[\frac{{^{n} C_0}^2}{n+1}+\frac{{^{n} C_1}^2}{n}+\frac{{^{n} C_2}^2}{n-1}+....\frac{{^{n} C_n}^2}{1}\]

OpenStudy (anonymous):

that notation is a little funky o.O are those: \[nCk = \left(\begin{matrix}n \\ k\end{matrix}\right) = \frac{n!}{k!(n-k)!}\] ?

hero (hero):

Hey dipankarstudy, you should check on that previous problem you posted the answer to jenmarhorne2.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and they are squared right? just trying to make sure im doing the right problem lol

OpenStudy (anonymous):

yes...that is an interesting problem ..i can give a hint...use (1+x)^n formula...

OpenStudy (anonymous):

no hints please =/ ive used that and gotten a formula without the squares, im still working on it though

OpenStudy (anonymous):

okay ...

OpenStudy (anonymous):

hello....got it?

OpenStudy (anonymous):

not yet =/ unfortunately.

OpenStudy (anonymous):

i have posted a new question u can try that also